# Rise of the forsaken relics: connecting present-day stellar streams and phase-mixed galaxies to the Epoch of Reionization

### v2

@@ Abstract
The ‘near-far’ approach to studying reionization leverages the star formation histories of the Milky Way (MW) or Local Group (LG) galaxies, derived from resolved photometry, to infer the low-mass/faint-end of the stellar mass functions (SMFs) or the ultraviolet luminosity functions (UVLFs) of high-redshift galaxies ($z\gtrsim 6$), beyond the current JWST detection limits ($M_{\mathrm{UV}}\gtrsim-15$). Previous works considered only intact low-mass galaxies in the MW and LG, neglecting disrupted galaxies such as stellar streams and phase-mixed objects. Using the FIRE-2 simulations, we show that these disrupted galaxies contribute up to $\sim$50% of the total stellar mass budget of the proto-MW/LG at z=69z=6-9. Including all the progenitors of these disrupted galaxies improves the normalization of the recovered SMFs/UVLFs by factors of $\sim$2–3 and reduces the halo-to-halo variation in the slope by $\sim$20–40%. This enables robust constraints down to at least the resolution limit of the simulations, near $M_{\star}\sim 10^{5}M_{\odot}$ or $M_{\mathrm{UV}}\sim-10$ at $z\gtrsim 6$. We also show that ‘fossil record’ reconstructions—which assume each present-day system descends from a single reionization-era progenitor—are sensitive to the stellar mass/UV magnitude thresholds, which introduces bias in the inferred low-mass/faint-end slopes. Additionally, we demonstrate that neglecting disrupted systems underestimates the contribution of galaxies with $M_{\mathrm{UV}}\lesssim-15$ to the reionization-era UV luminosity density. Finally, we estimate that a significant fraction ($\sim 50\%$) of streams with $M_{\star}\gtrsim 10^{6}M_{\odot}$ at zz=0 should be detectable from upcoming Rubin Observatory and Roman Space Telescope observations.
@@ I Introduction
However, some reionization models predict that fainter yet more numerous low-mass galaxies are major contributors to the total budget of ionizing photons, potentially with $M_{\mathrm{UV}}$ down to $\sim-8$ (Kuhlen & Faucher-Giguère, 2012; Wise et al., 2014; Robertson et al., 2015; Atek et al., 2015; Ma et al., 2018a, b; Finkelstein et al., 2019; Atek et al., 2024; Mascia et al., 2024; Simmonds et al., 2024; Harshan et al., 2024; Wu & Kravtsov, 2024). Many of these sources will remain beyond the reach of even JWST (Boylan-Kolchin et al., 2015, 2016; Weisz & Boylan-Kolchin, 2017). While observations of lensed galaxies can probe fainter magnitudes (down to $M_{\mathrm{UV}}\sim-12$), they introduce significant systematic uncertainties due to magnification effects (Atek et al., 2018; Bouwens et al., 2022a). Additionally, low-mass galaxies provide valuable insights into the impact of stellar feedback and the nature of dark matter on galaxy formation, particularly at high redshifts (Sun & Furlanetto, 2016; Bullock & Boylan-Kolchin, 2017; Sales et al., 2022; Sipple & Lidz, 2024; Sipple et al., 2025). Novel approaches for probing the faint/low-mass-end of the $z\gtrsim 6$ galaxy UV luminosity function (UVLF)/stellar mass function (SMF) are hence valuable not only in determining the role of low-luminosity galaxies in reionizing the universe, but also in understanding feedback processes and the properties of dark matter.
@@
One promising technique is the ‘near-far’ approach, which relies on studies of Milky Way (MW) satellite galaxies and Local Group (LG) low-mass galaxies at the present-day ($z=0$) to infer properties of their $z\gtrsim 6$ progenitors (Boylan-Kolchin et al., 2015, 2016; Weisz & Boylan-Kolchin, 2017; Weisz et al., 2014a). This employs color-magnitude diagrams (CMDs) of nearby MW/LG low-mass galaxies from resolved HST/JWST photometry to reconstruct their star-formation histories (SFHs) (Weisz et al., 2014b, a) and UV luminosities at $z\gtrsim 6$ (Weisz et al., 2014c; Weisz & Boylan-Kolchin, 2017), using stellar population synthesis models. This indirect approach probes fainter magnitudes than direct high-redshift observations, thereby providing information about the faint-end/low-mass-end of the UVLFs/SMFs at $z\gtrsim 6$.
@@
Previous works applied this technique to intact low-mass galaxies (Boylan-Kolchin et al., 2016; Weisz & Boylan-Kolchin, 2017; Weisz et al., 2014c; Skillman et al., 2017; Savino et al., 2023; McQuinn et al., 2024) in the LG. This strategy assumes a one-to-one mapping from high-$z$ low-mass galaxies to their present-day LG counterparts, ignoring the cosmological assembly histories of the low-mass galaxies and, in the case of satellite galaxies, their host halos. One consequence of this assumption is that the high-$z$ SMF/UVLF inferred from the LG fossil record recovers only $\approx$15-20% of the total amplitude (Gandhi et al., 2024). This is because it lacks information about the different progenitors that merged to form the present-day system. Gandhi et al. (2024) analyzed the impact of these mergers by including multiple progenitors of present-day low-mass galaxies in cosmological baryonic zoom-in simulations, and tested the near-far technique in terms of recovering the slope of the low-mass-end of the SMF at $z\gtrsim 6$. They found that the fossil record of surviving low-mass galaxies in the MW/LG can recover the slope of the faint-/low-mass-end of the UVLF/SMF of the proto-MW/LG at $z\gtrsim 6$, albeit with a large halo-to-halo scatter.
@@
Although these studies illustrate the general promise of the near-far technique, they do not yet fully account for the fact that most high-$z$ progenitors fail to survive intact as present-day MW/LG low-mass galaxies (Santistevan et al., 2020; Horta et al., 2024). In fact, most high-$z$ progenitor galaxies will be completely disrupted by mergers and tidal forces ($\sim 50\%$ of the proto-MW/LG, Gandhi et al. 2024), leaving behind both coherent stellar streams and galaxy remnants that are now part of the equilibrium phase-space distribution defined by the host galaxy (‘phase-mixed galaxies’). These merged substructures have systematically different properties from the intact satellites, such as lower metallicities due to earlier star formation (e.g. Panithanpaisal et al., 2021; Cunningham et al., 2022; Horta et al., 2023). The near-far technique usually ignores these disrupted galaxies when inferring the high-$z$ UVLF/SMF. However, the high-$z$ progenitors of these substructures also contribute to the ionizing photon budget at $z\gtrsim 6$, and their exclusion may lead to inaccurate conclusions regarding the role of low-luminosity galaxies in reionizing the universe. The reconstructed SMFs/UVLFs from intact galaxies alone require large corrections to the normalizations of these functions. Furthermore, simulations show a large scatter in the shape of the reconstructed SMFs/UVLFs across different MW/LG-like halos (halo-to-halo scatter) (Gandhi et al., 2024). Incorporating disrupted galaxies into the analysis could potentially help reduce this scatter and provide more robust constraints on the shape and normalization of reconstructed high-$z$ MW/LG SMF/UVLF. Hence, a comprehensive study of the impact of including disrupted galaxies on the near-far technique is crucial.
@@
In this paper, we investigate these issues using a suite of simulated MW/LG-like galaxies from zoomed cosmological hydrodynamic simulations. In section II, we describe the simulations used in this work. Section III discusses our criteria for classifying present-day substructures and our methodology for tracking them back to their simulated high-$z$ progenitors. Through Sections IV to VIII, we illustrate the impact of adding disrupted galaxies to the near-far technique and also test the feasibility of using them by imposing detectability constrains. Finally, we summarize these results in Section IX.
@@ II.2 Identifying low-mass galaxies
(× $\times\,10^{12}\,M_{\odot}$) (× $\times\,10^{10}\,M_{\odot}$) ($\rm{kpc}$) satellite galaxies stellar streams phase-mixed objects field galaxies
@@
Note. — Simulation name: References listed below; top 7 are isolated MW-like and bottom three are paired MW/M31-like. $M_{\rm 200m}$ and $R_{\rm 200m}$: total mass and spherical radius within which the mean density is 200 times the matter density of the universe. $M_{\star,90}$: stellar mass within a spherical radius that encloses 90% of the stellar mass within 20 kpc. $z_{\mathrm{MR_{3:1}}}$: the redshift at which the stellar mass ratio of the most massive halo to the second most massive subhalo is 3:1, which is when the dominant host galaxy emerges. Numbers of surviving satellite galaxies ($M_{\star}\gtrsim 10^{4.5}M_{\odot}$), stellar streams ($M_{\star}\gtrsim 10^{5.5}M_{\odot}$), phase-mixed galaxies ($M_{\star}\gtrsim 10^{5.5}M_{\odot}$), and field galaxies ($M_{\star}\gtrsim 10^{4.5}M_{\odot}$) are determined as described in Section III.1.
References. — A: Wetzel et al. (2016), B: Garrison-Kimmel et al. (2019a), C: Garrison-Kimmel et al. (2017), D: Hopkins et al. (2018a), E: Samuel et al. (2020), F: Garrison-Kimmel et al. (2019b).
III Present-day substructures and their high redshift progenitors
@@ IV Stellar mass distribution of progenitor galaxies at high redshifts
IV.1 Stellar mass budget during the EoR
@@ IV.2 Hierarchical structure formation at high redshifts
Gandhi et al. (2024) explored the effects of mergers and disruptions for present-day surviving low-mass galaxies by quantifying their progenitor counts during the reionization era. They found that the lowest mass simulated galaxies ($M_{\star}$ $\sim$ $10^{5}$ $M_{\odot}$) have $\sim$1–2 progenitors at $z\gtrsim 6$, while the progenitor number increases with present-day stellar mass, with the most massive galaxies ($M_{\star}$ $\sim$ $10^{9}$ $M_{\odot}$) having $\sim$15–40 progenitors. Furthermore, they showed that the fractional contribution of the main progenitor galaxy relative to the total stellar mass from all progenitor galaxies at that redshift, is $\gtrsim 50\%$ dominant across a wide range of present-day stellar masses ($M_{\star}$ $\sim$ 104.5-8 $M_{\odot}$). The present-day lowest-mass galaxies have an order unity contribution from a single main progenitor since they have at most $\sim$1–2 progenitors. On the other hand, high-mass present-day substructures are typically assembled from many smaller units, and the main progenitor makes a smaller fractional contribution to the total progenitor stellar mass at high redshift. Nevertheless, the main progenitor of even the most massive galaxies contributes $\sim$50% to the total progenitor stellar mass. Ultimately, Gandhi et al. (2024) concluded that the assumption of a single progenitor in case of the surviving low-mass galaxy fossil record still allows an accurate recovery of the low-mass/faint-end slope of the high-$z$ proto-MW/LG SMFs/UVLFs.
@@
V Stellar Mass Functions and UV Luminosity Functions at High-Redshifts
We further test the fossil record assumption by reconstructing the high-$z$ SMFs/UVLFs including disrupted galaxies. We examine how well the SMFs/UVLFs of the proto-MW/LG (and by extension the universal SMF/UVLF) at $z\gtrsim 6$ can be recovered from each present-day classification category, as well as from several different combinations of intact and disrupted galaxies. For now, our analysis assumes that the SFHs of these substructures can be precisely reconstructed from their CMDs at $z=0$. We compute the reconstructed SMFs/UVLFs in both the isolated and paired suite, and compare with the SMF/UVLF from FIREboxHR, which we treat as a benchmark for the universal SMF/UVLF. We do this for both the differential and cumulative functions.
@@ V Stellar Mass Functions and UV Luminosity Functions at High-Redshifts
These procedures result in reconstructed high-$z$ SMFs/UVLFs for both the all progenitors and fossil record scenarios. We compare them to universal benchmark SMFs/UVLFs in sections V.2 and V.3.
@@
V.1 Reconstructed differential SMFs and UVLFs for each present-day classification
@@
Figure 4 compares the differential SMFs (top panel) and UVLFs (bottom panel) of the proto-MW/LG (black) at $z=7$ with the ones constructed from the progenitors of present-day satellite galaxies (red), stellar streams (blue), phase-mixed galaxies (green), and field galaxies (yellow). The solid lines show reconstructions from the full set of progenitors and the dashed lines show reconstructions obtained using a simple model for the detectability of present-day substructures with the LSST (section VIII). We show results for the isolated suite and the paired suite in the left and right panels, respectively. The comparison between the two panels illustrates how well the reconstructed SMFs/UVLFs from each classification compare with the proto-MW/LG SMF/UVLF, when the near-far technique is applied to each of the MW and the LG alone. The curves show the median across all of the simulations in the respective galaxy suite. To highlight differences between classifications, the SMFs/UVLFs are computed in the all progenitors scenario rather than the fossil record scenario.
@@
The shape of the SMFs/UVLFs in the low-mass-/faint-end ($M_{\star}\lesssim 10^{6}M_{\odot}$ or $M_{UV}\gtrsim-13$) reconstructed from the progenitors of present-day stellar streams and phase-mixed galaxies is qualitatively similar in representing in the proto-MW/LG SMFs/UVLFs, when compared to the SMFs/UVLFs from just the progenitors of intact galaxies. Furthermore, the disrupted galaxy progenitors recovers more of the normalization of the SMFs/UVLFs at least in comparison to those from intact present-day satellite galaxies. This underlines the need to include present-day disrupted galaxies in near-far analysis. The SMFs/UVLFs from the disrupted galaxies tend to extend to more massive/brighter galaxies, because more of these systems have massive progenitors at high redshifts than in the intact galaxy population. The larger normalization of the SMFs/UVLFs from these disrupted systems, compared to that recovered from the low-mass satellites alone, is partly because some of these massive systems have multiple high-$z$ progenitors. The overall shapes do not change drastically even when detectability constraints are imposed, suggesting that it may be feasible to include disrupted galaxies in future near-far technique analyses. The SMFs and UVLFs reconstructed from the stream progenitors remain almost unchanged, because the detectable streams are the most massive ones with the largest number progenitors, comprising a large portion of the distribution at high redshifts.
@@
V.2 Reconstructed differential SMFs and UVLFs using the full merger histories of present-day substructures
@@
Figure 5: Differential SMFs at $z=7$ in the all progenitors scenario. In all the panels, red, blue, and green solid lines show the median across all the simulations for sample I, II, and III, respectively in the isolated (left) and paired (right) galaxy suites. The shaded regions show the corresponding 68% halo-to-halo scatter. In this scenario we measure all progenitors separately, instead of treating all progenitors as a single galaxy, as in the case of the fossil record. Top: reconstructed differential SMFs at $z=7$ from all of the progenitors in each present-day sample, compared to the proto-MW/LG (black; includes the present-day main host(s)) and the universal differential SMF from FIREboxHR (dashed yellow). Bottom: ratio of the SMF from each sample to the SMF of the proto-MW/LG. Reconstructed SMFs from sample II and III represent both the shape and normalization of the proto-MW/LG SMF better than sample I at all stellar masses, and generally appears to have smaller halo-to-halo scatter with some exceptions (see text for details).
@@
Figure 6: Differential UVLFs at $z=7$ in the all progenitors scenario. Same as Figure 5, but showing the differential UVLFs. As for the SMFs, including present-day streams (Sample II, blue) and phase-mixed galaxies (Sample III, green) improves the shape and normalization of the reconstructed UVLF compared to intact dwarf galaxies alone (Sample I, magenta).
@@
In this section, we explore how accurately we can recover the SMFs/UVLFs of the proto-MW/LG at $z\gtrsim 6$ by adding the progenitors of present-day disrupted galaxies along with the intact galaxies in the all progenitors scenario. Here we divide our present-day substructures into three samples:
@@
Figures 5 and 6 (top panels) compare the differential SMFs and UVLFs at $z=7$ for Samples I (red), II (blue), and III (green) to those of the proto-MW/LG (black) and the universal SMFs/UVLFs from FIREboxHR (yellow). As discussed in section V.1, we show the isolated suite (left) and the paired suite (right) separately. Solid lines show median SMFs/UVLFs from all of the simulations in the respective suite. Shaded regions show the 68% halo-to-halo scatter except for the universal FIREboxHR curve (yellow); here the shading denotes the 68% confidence interval (CI) obtained via bootstrapping (Feldmann et al., 2025). We also compare the ratio of the SMF/UVLF from each sample with respect to the SMF/UVLF of the proto-MW/LG (Figures 5 and 6, bottom panels) to help visualize their shapes as a function of stellar mass/UV magnitude. We compute the ratios for each simulation first and then take the median and 68% CI.

### v3

@@ Abstract
The “near-far” approach to studying reionization leverages the star formation histories of the Milky Way (MW) or Local Group (LG) galaxies, derived from resolved photometry, to infer the low-mass/faint end of the stellar mass functions (SMFs) or the ultraviolet luminosity functions (UVLFs) of high-redshift galaxies ($z\gtrsim 6$), beyond the current JWST detection limits ($M_{\mathrm{UV}}\gtrsim-15$). Previous works considered only intact low-mass galaxies in the MW and LG, neglecting disrupted galaxies such as stellar streams and phase-mixed objects. Using the FIRE-2 simulations, we show that these disrupted galaxies contribute up to $\sim$50% of the total stellar mass budget of the proto-MW/LG at z=69z=6-9. Including all the progenitors of these disrupted galaxies improves the normalization of the recovered SMFs/UVLFs by factors of $\sim$2–3 and reduces the halo-to-halo variation in the slope by $\sim$20–40%. This enables robust constraints down to at least the resolution limit of the simulations, near $M_{\star}\sim 10^{5}M_{\odot}$ or $M_{\mathrm{UV}}\sim-10$ at $z\gtrsim 6$. We also show that “fossil record” reconstructions—which assume each present-day system descends from a single reionization-era progenitor—are sensitive to the stellar-mass/UV-magnitude thresholds, which introduces bias in the inferred slopes at the low-mass/faint end. Additionally, we demonstrate that neglecting disrupted systems underestimates the contribution of galaxies with $M_{\mathrm{UV}}\lesssim-15$ to the reionization-era UV luminosity density. Finally, we estimate that a significant fraction ($\sim 50\%$) of streams with $M_{\star}\gtrsim 10^{6}M_{\odot}$ at zz=0 should be detectable from upcoming Rubin Observatory and Roman Space Telescope observations.
Galaxy evolution; Reionization; High-redshift galaxies; Stellar streams; Local Group; Milky Way Galaxy
@@ I Introduction
However, some reionization models predict that fainter yet more numerous low-mass galaxies are major contributors to the total budget of ionizing photons, potentially with $M_{\mathrm{UV}}$ down to $\sim-8$ (Kuhlen & Faucher-Giguère, 2012; Wise et al., 2014; Robertson et al., 2015; Atek et al., 2015; Ma et al., 2018a, b; Finkelstein et al., 2019; Atek et al., 2024; Mascia et al., 2024; Simmonds et al., 2024; Harshan et al., 2024; Wu & Kravtsov, 2024). Many of these sources will remain beyond the reach of even JWST (Boylan-Kolchin et al., 2015, 2016; Weisz & Boylan-Kolchin, 2017). While observations of lensed galaxies can probe fainter magnitudes (down to $M_{\mathrm{UV}}\sim-12$), they introduce significant systematic uncertainties due to magnification effects (Atek et al., 2018; Bouwens et al., 2022a). Additionally, low-mass galaxies provide valuable insights into the impact of stellar feedback and the nature of dark matter on galaxy formation, particularly at high redshifts (Sun & Furlanetto, 2016; Bullock & Boylan-Kolchin, 2017; Sales et al., 2022; Sipple & Lidz, 2024; Sipple et al., 2025). Novel approaches for probing the faint/low-mass end of the $z\gtrsim 6$ galaxy UV luminosity function (UVLF)/stellar mass function (SMF) are hence valuable not only in determining the role of low-luminosity galaxies in reionizing the universe, but also in understanding feedback processes and the properties of DM.
@@
One promising technique is the “near-far” approach, which relies on studies of Milky Way (MW) satellite galaxies and Local Group (LG) low-mass galaxies at the present-day ($z=0$) to infer properties of their $z\gtrsim 6$ progenitors (Boylan-Kolchin et al., 2015, 2016; Weisz & Boylan-Kolchin, 2017; Weisz et al., 2014a). This employs color-magnitude diagrams (CMDs) of nearby MW/LG low-mass galaxies from resolved HST/JWST photometry to reconstruct their star-formation histories (SFHs; Weisz et al. 2014b, a) and UV luminosities at $z\gtrsim 6$ (Weisz et al., 2014c; Weisz & Boylan-Kolchin, 2017), using stellar population synthesis models. This indirect approach probes fainter magnitudes than direct high-redshift observations, thereby providing information about the faint-end/low-mass-end of the UVLFs/SMFs at $z\gtrsim 6$.
@@
Previous works applied this technique to intact low-mass galaxies in the LG (Boylan-Kolchin et al., 2016; Weisz & Boylan-Kolchin, 2017; Weisz et al., 2014c; Skillman et al., 2017; Savino et al., 2023; McQuinn et al., 2024). This strategy assumes a one-to-one mapping from high-$z$ low-mass galaxies to their present-day LG counterparts, ignoring the cosmological assembly histories of the low-mass galaxies and, in the case of satellite galaxies, their host halos. One consequence of this assumption is that the high-$z$ SMF/UVLF inferred from the LG fossil record recovers only $\approx$15-20% of the total amplitude (Gandhi et al., 2024). This is because it lacks information about the different progenitors that merged to form the present-day system. Gandhi et al. (2024) analyzed the impact of these mergers by including multiple progenitors of present-day low-mass galaxies in cosmological baryonic zoom-in simulations, and tested the near-far technique in terms of recovering the slope of the low-mass-end of the SMF at $z\gtrsim 6$. They found that the fossil record of surviving low-mass galaxies in the MW/LG can recover the slope at the faint/low-mass end of the UVLF/SMF of the proto-MW/LG at $z\gtrsim 6$, albeit with a large halo-to-halo scatter.
@@
Although these studies illustrate the general promise of the near-far technique, they do not yet fully account for the fact that most high-$z$ progenitors fail to survive intact as present-day MW/LG low-mass galaxies (Santistevan et al., 2020; Horta et al., 2024). In fact, most high-$z$ progenitor galaxies will be completely disrupted by mergers and tidal forces ($\sim 50\%$ of the proto-MW/LG, Gandhi et al. 2024), leaving behind both coherent stellar streams and galaxy remnants that are now part of the equilibrium phase-space distribution defined by the host galaxy (“phase-mixed galaxies”). These merged substructures have systematically different properties from the intact satellites, such as lower metallicities due to earlier star formation (e.g., Panithanpaisal et al., 2021; Cunningham et al., 2022; Horta et al., 2023). The near-far technique usually ignores these disrupted galaxies when inferring the high-$z$ UVLF/SMF. However, the high-$z$ progenitors of these substructures also contribute to the ionizing photon budget at $z\gtrsim 6$, and their exclusion may lead to inaccurate conclusions regarding the role of low-luminosity galaxies in reionizing the universe. The reconstructed SMFs/UVLFs from intact galaxies alone require large corrections to the normalizations of these functions. Furthermore, simulations show a large scatter in the shape of the reconstructed SMFs/UVLFs across different MW/LG-like halos (halo-to-halo scatter) (Gandhi et al., 2024). Incorporating disrupted galaxies into the analysis could potentially help reduce this scatter and provide more robust constraints on the shape and normalization of reconstructed high-$z$ MW/LG SMF/UVLF. Hence, a comprehensive study of the impact of including disrupted galaxies on the near-far technique is crucial.
@@
In this paper, we investigate these issues using a suite of simulated MW/LG-like galaxies from zoomed cosmological hydrodynamic simulations. In Section II, we describe the simulations used in this work. Section III discusses our criteria for classifying present-day substructures and our methodology for tracking them back to their simulated high-$z$ progenitors. Through Sections IVVIII, we illustrate the impact of adding disrupted galaxies to the near-far technique and also test the feasibility of using them by imposing detectability constrains. Finally, we summarize these results in Section IX.
@@ II.2 Identifying Low-mass Galaxies
($\times 10^{12}\,M_{\odot}$) ($\times 10^{10}\,M_{\odot}$) ($\rm{kpc}$) Satellite Galaxies Stellar Streams Phase-mixed Objects Field Galaxies
@@
Note. — Simulation Name: References listed below; the top seven are isolated MW-like and bottom three are paired MW/M31-like. $M_{\rm 200m}$ and $R_{\rm 200m}$: total mass and spherical radius within which the mean density is 200 times the matter density of the Universe. $M_{\star,90}$: stellar mass within a spherical radius that encloses 90% of the stellar mass within 20 kpc. $z_{\mathrm{MR_{3:1}}}$: the redshift at which the stellar-mass ratio of the most massive halo to the second most massive subhalo is 3:1, which is when the dominant host galaxy emerges. Numbers of surviving satellite galaxies ($M_{\star}\gtrsim 10^{4.5}M_{\odot}$), stellar streams ($M_{\star}\gtrsim 10^{5.5}M_{\odot}$), phase-mixed galaxies ($M_{\star}\gtrsim 10^{5.5}M_{\odot}$), and field galaxies ($M_{\star}\gtrsim 10^{4.5}M_{\odot}$) are determined as described in Section III.1.
References. — (A) Wetzel et al. (2016), (B) Garrison-Kimmel et al. (2019a), (C) Garrison-Kimmel et al. (2017), (D) Hopkins et al. (2018a), (E) Samuel et al. (2020), (F) Garrison-Kimmel et al. (2019b).
III Present-day Substructures and their High-redshift Progenitors
@@ IV Stellar mass distribution of progenitor galaxies at high redshifts
IV.1 Stellar-mass Budget During the Epoch of Reionization
@@ IV.2 Hierarchical Structure Formation at High Redshifts
Gandhi et al. (2024) explored the effects of mergers and disruptions for present-day surviving low-mass galaxies by quantifying their progenitor counts during the reionization era. They found that the lowest mass simulated galaxies ($M_{\star}$ $\sim$ $10^{5}$ $M_{\odot}$) have $\sim$1–2 progenitors at $z\gtrsim 6$, while the progenitor number increases with present-day stellar mass, with the most massive galaxies ($M_{\star}$ $\sim$ $10^{9}$ $M_{\odot}$) having $\sim$15–40 progenitors. Furthermore, they showed that the fractional contribution of the main progenitor galaxy relative to the total stellar mass from all progenitor galaxies at that redshift, is $\gtrsim 50\%$ dominant across a wide range of present-day stellar masses ($M_{\star}$ $\sim$ 104.5-8 $M_{\odot}$). The present-day lowest-mass galaxies have an order unity contribution from a single main progenitor since they have at most $\sim$1–2 progenitors. On the other hand, high-mass present-day substructures are typically assembled from many smaller units, and the main progenitor makes a smaller fractional contribution to the total progenitor stellar mass at high redshift. Nevertheless, the main progenitor of even the most massive galaxies contributes $\sim$50% to the total progenitor stellar mass. Ultimately, Gandhi et al. (2024) concluded that the assumption of a single progenitor in case of the surviving low-mass galaxy fossil record still allows an accurate recovery of the slopes at the low-mass/faint end of the high-$z$ proto-MW/LG SMFs/UVLFs.
@@
V Stellar Mass Functions and Ultraviolet Luminosity Functions at High-Redshifts
We further test the fossil-record assumption by reconstructing the high-$z$ SMFs/UVLFs including disrupted galaxies. We examine how well the SMFs/UVLFs of the proto-MW/LG (and by extension the universal SMF/UVLF) at $z\gtrsim 6$ can be recovered from each present-day classification category, as well as from several different combinations of intact and disrupted galaxies. For now, our analysis assumes that the SFHs of these substructures can be precisely reconstructed from their CMDs at $z=0$. We compute the reconstructed SMFs/UVLFs in both the isolated and paired suite, and compare with the SMF/UVLF from FIREboxHR, which we treat as a benchmark for the universal SMF/UVLF. We do this for both the differential and cumulative functions.
@@ V Stellar Mass Functions and Ultraviolet Luminosity Functions at High-Redshifts
These procedures result in reconstructed high-$z$ SMFs/UVLFs for both the all-progenitors and fossil-record scenarios. We compare them to universal benchmark SMFs/UVLFs in sections V.2 and V.3.
@@
V.1 Reconstructed Differential Stellar Mass Functions and Ultraviolet Luminosity Functions for Each Present-day Classification
@@
Figure 4 compares the differential SMFs (top panel) and UVLFs (bottom panel) of the proto-MW/LG (black) at $z=7$ with the ones constructed from the progenitors of present-day satellite galaxies (red), stellar streams (blue), phase-mixed galaxies (green), and field galaxies (yellow). The solid lines show reconstructions from the full set of progenitors and the dashed lines show reconstructions obtained using a simple model for the detectability of present-day substructures with the LSST (section VIII). We show results for the isolated suite and the paired suite in the left and right panels, respectively. The comparison between the two panels illustrates how well the reconstructed SMFs/UVLFs from each classification compare with the proto-MW/LG SMF/UVLF, when the near-far technique is applied to each of the MW and the LG alone. The curves show the median across all of the simulations in the respective galaxy suite. To highlight differences between classifications, the SMFs/UVLFs are computed in the all-progenitors scenario rather than the fossil-record scenario.
@@
The shape of the SMFs/UVLFs in the low-mass/faint end ($M_{\star}\lesssim 10^{6}M_{\odot}$ or $M_{UV}\gtrsim-13$) reconstructed from the progenitors of present-day stellar streams and phase-mixed galaxies is qualitatively similar in representing in the proto-MW/LG SMFs/UVLFs, when compared to the SMFs/UVLFs from just the progenitors of intact galaxies. Furthermore, the disrupted galaxy progenitors recovers more of the normalization of the SMFs/UVLFs at least in comparison to those from intact present-day satellite galaxies. This underlines the need to include present-day disrupted galaxies in near-far analysis. The SMFs/UVLFs from the disrupted galaxies tend to extend to more massive/brighter galaxies, because more of these systems have massive progenitors at high redshifts than in the intact galaxy population. The larger normalization of the SMFs/UVLFs from these disrupted systems, compared to that recovered from the low-mass satellites alone, is partly because some of these massive systems have multiple high-$z$ progenitors. The overall shapes do not change drastically even when detectability constraints are imposed, suggesting that it may be feasible to include disrupted galaxies in future near-far technique analyses. The SMFs and UVLFs reconstructed from the stream progenitors remain almost unchanged, because the detectable streams are the most massive ones with the largest number progenitors, comprising a large portion of the distribution at high redshifts.
@@
V.2 Reconstructed Differential Stellar Mass Functions and Ultraviolet Luminosity Functions Using the Full Merger Histories of Present-day Substructures
@@
Figure 5: Differential SMFs at $z=7$ in the all-progenitors scenario. In all the panels, red, blue, and green solid lines show the median across all the simulations for Samples I, II, and III, respectively in the isolated (left) and paired (right) galaxy suites. The shaded regions show the corresponding 68% halo-to-halo scatter. In this scenario we measure all progenitors separately, instead of treating all progenitors as a single galaxy, as in the case of the fossil record. Top: reconstructed differential SMFs at $z=7$ from all of the progenitors in each present-day sample, compared to the proto-MW/LG (black; includes the present-day main host(s)) and the universal differential SMF from FIREboxHR (dashed yellow). Bottom: ratio of the SMF from each sample to the SMF of the proto-MW/LG. Reconstructed SMFs from sample II and III represent both the shape and normalization of the proto-MW/LG SMF better than sample I at all stellar masses, and generally appears to have smaller halo-to-halo scatter with some exceptions (see text for details).
@@
Figure 6: Differential UVLFs at $z=7$ in the all-progenitors scenario. Same as Figure 5, but showing the differential UVLFs. As for the SMFs, including present-day streams (Sample II, blue) and phase-mixed galaxies (Sample III, green) improves the shape and normalization of the reconstructed UVLF compared to intact dwarf galaxies alone (Sample I, magenta).
@@
In this section, we explore how accurately we can recover the SMFs/UVLFs of the proto-MW/LG at $z\gtrsim 6$ by adding the progenitors of present-day disrupted galaxies along with the intact galaxies in the all-progenitors scenario. Here we divide our present-day substructures into three samples:
@@
Figures 5 and 6 (top panels) compare the differential SMFs and UVLFs at $z=7$ for Samples I (red), II (blue), and III (green) to those of the proto-MW/LG (black) and the universal SMFs/UVLFs from FIREboxHR (yellow). As discussed in section V.1, we show the isolated suite (left) and the paired suite (right) separately. Solid lines show median SMFs/UVLFs from all of the simulations in the respective suite. Shaded regions show the 68% halo-to-halo scatter except for the universal FIREboxHR curve (yellow); here the shading denotes the 68% confidence interval (CI) obtained via bootstrapping (Feldmann et al., 2025). We also compare the ratio of the SMF/UVLF from each sample with respect to the SMF/UVLF of the proto-MW/LG (Figures 5 and 6, bottom panels) to help visualize their shapes as a function of stellar-mass/UV-magnitude. We compute the ratios for each simulation first and then take the median and 68% CI.